OpenStudy (anonymous):

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length. Help!

OpenStudy (jack1):



OpenStudy (jack1):

\[h^2 = a^2 + b^2 \]

OpenStudy (jack1):

so if a = 1 and b = 1... h^2 = ...? therefore h = ...?

OpenStudy (jack1):

@yulisasandoval15 ?

OpenStudy (anonymous):

ok i get this now.. thanks...

OpenStudy (jack1):

sweet, yw

OpenStudy (anonymous):

yeah